OpenStudy (anonymous):

I want to build a right triangular deck behind my house. The hypotenuse of the deck will be 40 feet long. The other two sides of the deck will be equal in length. What is the length of the deck?Answer 56.57 feet 6.32 feet 4.47 feet 28.28 feet

OpenStudy (anonymous):

@Jithinkv

OpenStudy (anonymous):

@Jithinkv

OpenStudy (abb0t):

\[a^2 + b^2 = c^2\]

OpenStudy (abb0t):

and \[a^2 = b^2\]

OpenStudy (anonymous):

2a^2 = c^2

OpenStudy (anonymous):

comes around 28.28